OpenStudy (anonymous):

how to solve this partial fraction: f(x) = 3x/(1+2x)^m where m is integer 1,2,....,m

OpenStudy (anonymous):

For m =8 \[ \frac{3 x}{(2 x+1)^8}=\frac{3}{2 (2 x+1)^7}-\frac{3}{2 (2 x+1)^8} \]

OpenStudy (anonymous):

for m=9 \[ \frac{3 x}{(2 x+1)^9}=\frac{3}{2 (2 x+1)^8}-\frac{3}{2 (2 x+1)^9} \]

OpenStudy (anonymous):

Can you find a pattern?

OpenStudy (anonymous):

Hi first thank you very much for the answer. but this is not the case that I need, in my case m is unknown, and then I can't said m=8 or any other number. the soultion of the bove equation is: \[\frac{ 3x }{(1+2x)^m } = \sum_{i=1}^{m} \frac{ Ai }{ (1+2x)^i }\] now I don't know how to find the value of Ai , that is my problem thank you

OpenStudy (anonymous):

All of them are zero except \[ A_m=-\frac 3 2 \text { and } A_{m-1=}\frac 3 2 \]